# Asynchronous Federated Stochastic Optimization with Exact Averaging for Heterogeneous Local Objectives

### v1

@@ 4 Convergence analysis
In this section, we present our theoretical results on the convergence of AREA. We first set up our probabilistic framework to derive the iterates of AREA and list our assumptions, then conclude with our theorems. The proofs for all theorems and intermediate lemmas can be found in Appendix A.
4.1 Setup
To more accurately represent the inconsistent behavior of clients in the federated setting, we have adopted a probabilistic setup that allows for arbitrarily long delays between updates. Specifically, we define an iteration of AREA to be an occurrence of one of the events $E_{i}$ for $i\in[n]$ and $E_{s}$, defined as follows: $E_{i}\triangleq\{\text{The server receives a request by client $i$}\}$ for $i\in[n]$, and $E_{s}\triangleq\{\text{The aggregation criterion is satisfied}\}$. We assume that the events $E_{s}$ and $E_{i}$ for $i\in[n]$ follow independent stochastic processes, i.e., the time elapsed between two updates from client $i$ is independent from the time elapsed between two updates from client $j\in[n]-\{i\}$ and from the time elapsed between two global aggregations.111 This assumption does not cover aggregation schemes where the receipt of a fixed number of client updates triggers a global update (e.g., [33]); however, since our theoretical results are derived under the assumption of no information on the arrivals of client updates, they can be interpreted as a worst case/upper bounds for aggregation strategies making use of this information. Moreover, let $J_{k}\in\{1,2,...,n,s\}$ be a random variable describing the index $j$ of the event $E_{j}$ occurring at iteration $k$. We define $p_{i}\triangleq\mathbb{P}(J_{k}=i)>0$ for $i\in[n]$ and $p_{s}\triangleq\mathbb{P}(J_{k}=s)>0$.
@@ 4.1 Setup
A simple instance accommodated by our setup involves modeling the time $t_{i}>0$ between two updates from client $i\in[n]$ and the time $t_{s}>0$ between two global aggregations at the server as independent but not identically distributed exponential random variables, i.e., $t_{i}\sim\exp(\lambda_{i})$ for $i\in[n]$ and $t_{s}\sim\exp(\lambda_{s})$. Then $\mathbb{P}(J_{k}=i)=\mathbb{P}(\min\{t_{1},...,t_{n},t_{s}\}=t_{i})=\lambda_{i}/\bar{\lambda}$ for $i\in[n]$ and $\mathbb{P}(J_{k}=s)=\mathbb{P}(\min\{t_{1},...,t_{n},t_{s}\}=t_{s})=\lambda_{s}/\bar{\lambda}$, where $\bar{\lambda}=\lambda_{s}+\sum_{i=1}^{n}\lambda_{i}$. This framework, known as Poisson clocks, is widely used to model asynchronous communication over networks [e.g., 42, 27].
Unlike the majority of concurrent works, we do not impose deterministic bounds on the maximum or average client delays, and allow for heterogeneous update rates among clients. Moreover, modeling the satisfaction of the aggregation criterion as a random variable permits the adoption of a wide variety of stochastic and deterministic aggregation policies. [^2] We note that the deterministic aggregation policies could be considered as special cases of the stochastic ones.
For each iteration $k\in\mathbb{N}$, we define the Bernoulli random variable $\phi_{s,k}$ (resp. $\phi_{i,k}$) taking the value $1$ if $J_{k}=s$ (resp. $J_{k}=i$) and the value 00 otherwise. It is easy to check that $\mathbb{E}[\phi_{s,k}|\mathcal{F}_{k-1}]=p_{s}$ and $\mathbb{E}[\phi_{i,k}|\mathcal{F}_{k-1}]=p_{i}$ for $i\in[n]$ and $k\in\mathbb{N}$. We can now express the iterates of AREA for $i\in[n]$ and $k=0,1,2,...$ as
$x_{i,k+1}$ $=x_{i,k}+\phi_{i,k+1}\left(P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-x_{i,k}\right),$
$y_{i,k+1}$ $=y_{i,k}+\phi_{i,k+1}(x_{i,k}-y_{i,k}),$
$u_{s,k+1}$ $=\sum_{i=1}^{n}\phi_{i,k+1}\left(u_{s,k}+\frac{1}{n}(x_{i,k}-y_{i,k})\right),$
@@
Note that at each iteration count $k=0,1,2,...$ we have $\phi_{s,k+1}+\sum_{i=1}^{n}\phi_{i,k+1}=1$ and $\phi_{s,k+1}\cdot\prod_{i=1}^{n}\phi_{i,k+1}=0$ by construction (i.e., only one element in the set $\{\phi_{1,k+1},...,\phi_{n,k+1},\phi_{s,k+1}\}$ is equal to $1$ and the remaining elements are 00). To further simplify the above iterates, let us define the variables $z_{i,k}\triangleq x_{s,k}+u_{s,k}-y_{i,k}$ for $i\in[n]$, and recall that $u_{s,0}=0$ and $y_{i,0}=x_{s,0}$ for $i\in[n]$. We then have $\bar{z}_{k+1}=x_{s,k+1}+u_{k+1}-\bar{y}_{k+1}=x_{s,k}+\left(\phi_{s,k+1}+\sum_{i=1}^{n}\phi_{i,k+1}\right)u_{s,k}-\bar{y}_{k}=\bar{z}_{k}$. Hence, $\bar{z}_{k}=\bar{z}_{0}=0$ for all $k\in\mathbb{N}$, and we can eliminate the aggregator variable by setting $u_{s,k}=\bar{y}_{k}-x_{s,k}$. The final form of the iterates of AREA then becomes
$x_{i,k+1}$ $=x_{i,k}+\phi_{i,k+1}\left(P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-x_{i,k}\right),$ (2)
$y_{i,k+1}$ $=y_{i,k}+\phi_{i,k+1}(x_{i,k}-y_{i,k}),$ (3)
@@ 4.2 Main results
We present our theoretical results on the convergence of AREA under two instances of Problem (1): (i) smooth and strongly convex $f_{i}(\cdot)$ for $i\in[n]$ over $\mathcal{X}=\mathbb{R}^{p}$ and (ii) convex nonsmooth $f_{i}(\cdot)$ for $i\in[n]$ over convex and compact $\mathcal{X}\subset\mathbb{R}^{p}$. Although these assumptions cover a limited subset of problems, they allow for a rigorous convergence analysis, providing valuable insights into the behavior of AREA. We first list a number of not necessarily overlapping assumptions on the functions $f_{i}(\cdot)$ of Problem (1).
@@
Next, we state our main result on the convergence of AREA under the assumptions of strong convexity and smoothness.
@@
Under Assumptions 5 and 6, suppose that ($i$) $\mathcal{X}=\mathbb{R}^{p}$ in Problem (1) and ($ii$) $\alpha_{k}=\alpha\in\left(0,\min_{i}\left\{\frac{2}{\mu_{i}+L_{i}}\right\}\right)$ in Eq. (2) for all $k\in\mathbb{N}_{0}$. Then AREA converges in the mean square sense to a neighborhood of the optimal solution of Problem (1) with linear rate $\mathcal{O}(\rho(A))<1$, where $\rho(A)$ is the spectral radius of the block matrix $A\in\mathbb{R}^{(2n+1)\times(2n+1)}$ given by
$A=\begin{bmatrix}I-P&0&P(1_{n}-\alpha\gamma)\\ P&I-P&0\\ 0&\frac{p_{s}}{n}1_{n}^{\prime}&1-p_{s}\end{bmatrix},$
@@
where $P=diag(p_{1},...,p_{n})\in\mathbb{R}^{n\times n}$, $1_{n}$ is the vector of ones of size $n$, $\gamma=[\gamma_{1},...,\gamma_{n}]^{\prime}\in\mathbb{R}^{n}$, and $\gamma_{i}=\frac{2\mu_{i}L_{i}}{\mu_{i}+L_{i}}$.
Moreover, the error neighborhood for the server iterates of AREA satisfies $\limsup_{k\rightarrow\infty}\mathbb{E}[\|x_{s,k}-x^{\star}\|^{2}]\leq\frac{\alpha\sigma^{2}}{\bar{\gamma}},$ where $\sigma^{2}$ is defined in Assumption 1 and $\bar{\gamma}=\frac{1}{n}\sum_{i=1}^{n}\gamma_{i}$.
Theorem 8 indicates that both the acceptable stepsize range and the size of the error neighborhood for AREA are independent of the probabilities $p_{s}$ and $p_{i}$ for $i\in[n]$; this contrasts with existing works on this topic that require prior knowledge of the maximum or average client delays to tune the stepsize. Our analysis also characterizes the dependency of the convergence rate of AREA on the probabilities $p_{s}$ and $p_{i}$ for $i\in[n]$ through the block matrix $A$. We note that exact convergence is achieved without the stochastic noise ($\sigma=0$ in Assumption 1).
@@
In the setting of Remark 4, if the client update frequencies $\lambda_{i}$ for $i\in[n]$ are known, one can numerically solve for the aggregation frequency $\lambda_{s}$, which maximizes the convergence rate (i.e., minimizes $\rho(A)$).
We conclude this section with a result on the convergence of AREA when the functions $f_{i}(\cdot)$ of Problem (1) are convex and nonsmooth. We note that in this version of AREA, clients have access only to stochastic approximations of the subgradients of their local functions. Hence, even in the absence of stochastic noise, a diminishing stepsize sequence is required to achieve convergence to the set of optimal solutions of Problem (1).
@@
In addition, the sequence of server iterates $\{x_{k}\}$ converges to a solution of Problem (1) with probability $1$.
Theorem 10 establishes the almost sure convergence of the iterates of AREA to an optimal solution of Problem (1) provided the stepsize sequence $\{\alpha_{k}\}$ is square summable but not summable. Moreover, taking the total expectation on both sides of Eq. (5) and applying recursively yields
$\sum_{t=0}^{K-1}\alpha_{k}\mathbb{E}[f(x_{s,k})-f^{\star}]\leq\frac{n}{2}\left(V_{0}+(\sigma^{2}+C^{2})\sum_{t=0}^{K-1}\alpha_{k}^{2}\right),$
and we recover the convergence rate of the centralized stochastic projected subgradient method [7].

### v2

@@ 4 Convergence analysis
We now present our theoretical results on the convergence of AREA. We first set up our probabilistic framework for deriving the iterates of AREA and list our assumptions, then conclude with our theorems. The proofs for all theorems and intermediate lemmas can be found in Appendix A.1.
To capture more accurately the inconsistent behavior of clients in the FL setting, we adopt a probabilistic setup that allows for arbitrarily long delays between updates. Specifically, we define as an iteration of AREA every occurrence of one of the events $E_{i}$ for $i\in[n]$ and $E_{s}$, defined as follows: $E_{i}\triangleq\{\text{The server receives a request by client $i$}\}$ for $i\in[n]$, and $E_{s}\triangleq\{\text{The aggregation criterion is satisfied}\}$. We assume that the events $E_{s}$ and $E_{i}$ for $i\in[n]$ follow independent stochastic processes, namely, the time elapsed between two consecutive updates from client $i$ is independent from the time between two updates from client $j\neq i$ and from the time between two aggregations. [^1] This assumption does not cover aggregation schemes where the receipt of a fixed number of client updates triggers a global update (e.g., [35]); however, since our theoretical results are derived under the assumption of no information on the arrivals of client updates, they can be interpreted as a worst case/upper bounds for aggregation strategies making use of this information. Let $J_{k}\in\{1,2,...,n,s\}$ be a random variable describing the index $j$ of the event $E_{j}$ occurring at iteration $k$. We define $p_{i}\triangleq\mathbb{P}(J_{k}=i)>0$ for $i\in[n]$ and $p_{s}\triangleq\mathbb{P}(J_{k}=s)>0$.
@@
A simple instance accommodated by our setup is modeling the time $t_{i}>0$ between two updates from client $i\in[n]$ and the time $t_{s}>0$ between two aggregations as independent exponential random variables with different parameters, i.e., $t_{i}\sim\exp(\lambda_{i})$ for $i\in[n]$ and $t_{s}\sim\exp(\lambda_{s})$. Then $\mathbb{P}(J_{k}=j)=\mathbb{P}(\min\{t_{1},...,t_{n},t_{s}\}=t_{j})$ which yields $p_{i}=\lambda_{i}/\bar{\lambda}$ for $i\in[n]$ and $p_{s}=\lambda_{s}/\bar{\lambda}$, where $\bar{\lambda}=\lambda_{s}+\sum_{i=1}^{n}\lambda_{i}$. This framework, sometimes referred to as Poisson clocks, is widely used to model asynchronous communication over networks, e.g., [45, 28, 9].
Unlike concurrent works, we do not impose deterministic bounds on the maximum or average client delays, or the growth of the delay sequence, and allow for heterogeneous client update frequencies. Moreover, modeling the satisfaction of the aggregation criterion as a random variable permits the adoption of a wide variety of stochastic and deterministic aggregation policies. [^2] Deterministic aggregation policies can be considered as special cases of the stochastic ones.
Let us define the Bernoulli random variables $\phi_{s,k}$ (resp. $\phi_{i,k}$) taking the value $1$ if $J_{k}=s$ (resp. $J_{k}=i$) at the $k^{th}$ iteration and 00 otherwise; then $\mathbb{E}[\phi_{s,k}|\mathcal{F}_{k-1}]=p_{s}$ and $\mathbb{E}[\phi_{i,k}|\mathcal{F}_{k-1}]=p_{i}$ for $i\in[n]$ and $k\in\mathbb{N}$. Moreover, assuming that the server keeps a count $k\in\mathbb{N}_{0}$ of the events $E_{i}$ for $i\in[n]$ and $E_{s}$, define $\alpha_{k}=h(\alpha_{k-1})$ for $k\in\mathbb{N}$ (we note that this counter is not necessary in practice due to the use of constant stepsizes; we only employ it for theoretical purposes). The iterates of AREA for $k\in\mathbb{N}_{0}$ can be expressed as
$x_{i,k+1}$ $=x_{i,k}+\phi_{i,k+1}\left(P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-x_{i,k}\right),\quad\forall i\in[n],$
$y_{i,k+1}$ $=y_{i,k}+\phi_{i,k+1}(x_{i,k}-y_{i,k}),\quad\forall i\in[n],$
$u_{s,k+1}$ $=\textstyle\sum_{i=1}^{n}\phi_{i,k+1}\left(u_{s,k}+\frac{1}{n}(x_{i,k}-y_{i,k})\right),$
@@
We note that by construction, since each iteration count $k\in\mathbb{N}$ corresponds to a single event from the set $\{E_{1},...,E_{n},E_{s}\}$, only a single variable in the set $\{\phi_{1,k},...,\phi_{n,k},\phi_{s,k}\}$ is equal to $1$ and the rest are 00. To further simplify the above iterates, let us define the variables $z_{i,k}\triangleq x_{s,k}+u_{s,k}-y_{i,k}$ for $i\in[n]$, and recall that $u_{s,0}=0$ and $y_{i,0}=x_{s,0}$ for $i\in[n]$. We then have $\bar{z}_{k+1}=x_{s,k+1}+u_{k+1}-\bar{y}_{k+1}=x_{s,k}+\left(\phi_{s,k+1}+\sum_{i=1}^{n}\phi_{i,k+1}\right)u_{s,k}-\bar{y}_{k}=\bar{z}_{k}$. Hence, $\bar{z}_{k}=\bar{z}_{0}=0$ for all $k\in\mathbb{N}$, and we can eliminate the aggregator variable by setting $u_{s,k}=\bar{y}_{k}-x_{s,k}$. The final form of the iterates of AREA then becomes
$x_{i,k+1}$ $=x_{i,k}+\phi_{i,k+1}\left(P_{\mathcal{X}}[x_{s,k}-\alpha_{k}g_{i}(x_{s,k})]-x_{i,k}\right),\quad\forall i\in[n],$ (2)
$y_{i,k+1}$ $=y_{i,k}+\phi_{i,k+1}(x_{i,k}-y_{i,k}),\quad\forall i\in[n],$ (3)
@@ 4.1 Main results
We provide theoretical results on the convergence of AREA for two instances of Problem (1): (i) smooth and strongly convex $f_{i}(\cdot)$ for $i\in[n]$ over $\mathcal{X}=\mathbb{R}^{p}$ and (ii) convex nonsmooth $f_{i}(\cdot)$ for $i\in[n]$ over convex and compact $\mathcal{X}\subset\mathbb{R}^{p}$. We first list a number of not necessarily overlapping assumptions on the functions $f_{i}(\cdot)$ of Problem (1).
@@
Next, we state our main result on the convergence of AREA for strongly convex, smooth functions.
@@
Under Assumptions 5 and 6, suppose that ($i$) $\mathcal{X}=\mathbb{R}^{p}$ in Problem (1) and ($ii$) $\alpha_{k}=\alpha\in\left(0,\min_{i}\left\{\frac{2}{\mu_{i}+L_{i}}\right\}\right)$ in Eq. (2) for all $k\in\mathbb{N}_{0}$. Moreover, define the points $x_{i}^{\star}\triangleq x^{\star}-\alpha\nabla f_{i}(x^{\star})\in\mathbb{R}^{p}$ for $i\in[n]$, where $x^{\star}$ is the optimal solution of Problem (1), and the vector $U_{k}\in\mathbb{R}^{2n+1}$ for $k\in\mathbb{N}_{0}$ where $[U_{k}]_{i}\triangleq\mathbb{E}[\|x_{i,k}-x_{i}^{\star}\|^{2}]$ and $[U_{k}]_{n+i}\triangleq\mathbb{E}[\|y_{i,k}-x_{i}^{\star}\|^{2}]$ for $i\in[n]$, and $[U_{k}]_{2n+1}\triangleq\mathbb{E}[\|x_{s,k}-x^{\star}\|^{2}]$. Then for $k\in\mathbb{N}_{0}$, the sequence $\{U_{k}\}$ satisfies the relation
$U_{k}\leq A^{k}U_{0}+\alpha^{2}\sigma^{2}\sum_{t=0}^{k-1}A^{t}q,$ (5)
@@
where $A=\begin{bmatrix}I-P&0&P(1_{n}-\alpha\gamma)\\ P&I-P&0\\ 0&\frac{p_{s}}{n}1_{n}^{\prime}&1-p_{s}\end{bmatrix}\in\mathbb{R}^{(2n+1)\times(2n+1)},\quad q=\begin{bmatrix}P1_{n}\\ 0\\ 0\end{bmatrix}\in\mathbb{R}^{2n+1}$, $P=diag(p_{1},...,p_{n})\in\mathbb{R}^{n\times n}$, $1_{n}$ is the vector of ones of size $n$, $\gamma=[\gamma_{1},...,\gamma_{n}]^{\prime}\in\mathbb{R}^{n}$, $\gamma_{i}=\frac{2\mu_{i}L_{i}}{\mu_{i}+L_{i}}$, and $\sigma^{2}$ is defined in Assumption 1.
Moreover, $\rho(A)<1$, and $\left[\sum_{t=0}^{k-1}A^{t}q\right]_{n+1}\leq\frac{1}{\alpha\bar{\gamma}}$ for $k\in\mathbb{N}_{0}$, where $\bar{\gamma}=\frac{1}{n}\sum_{i=1}^{n}\gamma_{i}$.
The result of Theorem 8 can be interpreted as follows: the expected distance to optimality of the server iterates of AREA at the $k^{th}$ iteration, i.e. $[U_{k}]_{2n+1}$, is characterized by the $(2n+1)^{th}$ rows of the two terms on the right-hand side of (5). The first term, $[A^{k}U_{0}]_{2n+1}$, decays exponentially fast with $k$ with rate $\mathcal{O}(\rho(A))$, while the second term is upper bounded by $\frac{\alpha\sigma^{2}}{\bar{\gamma}}$ for $k\in\mathbb{N}_{0}$. We note that the second term is independent of the probabilities $p_{i}$ for $i\in[n]$ and $p_{s}$, which appear only in the matrix $A$, and hence influence $\rho(A)$. Although it is not possible to find a closed form expression for $\rho(A)$ and obtain tight convergence rates for AREA, when the second term dominates the right-hand side of (5), we obtain error $\mathcal{O}\left(\sigma^{2}/T\right)$ after $T$ iterations with $\alpha=\mathcal{O}(\bar{\gamma}/T)$. This bound is independent of the heterogeneity between the local functions, theoretically validating the effectiveness of our mechanism in eliminating the client drift due to asynchrony.
To the best of our knowledge, convergence guarantees under non-iid data and strong convexity and smoothness are provided only for asynchronous SGD in [8], and the methods AsyncCommSGD [3] and MIFA [17], that employ error-correction schemes similar to ours. The work [8] obtains a convergence rate of $\mathcal{O}\left(\left(n\sigma^{2}+\zeta^{2}\sum_{i=1}^{n}\tau_{i}^{2}\right)/T\right)$, where $\tau_{i}$ is the delay of client $i$ and $\zeta=\max_{i}\|\nabla f_{i}(x)-\nabla f(x)\|$; the last term grows with the heterogeneity between the local datasets and has a quadratic dependency on client delays, suggesting that staleness exacerbates client drift for asynchronous SGD. Conversely, [3, 17] obtain a rate of $\mathcal{O}\left(\sigma^{2}/(nT)+\tau_{\max}^{2}D^{2}/T^{2}\right)$ for the sequence $\{f(\bar{x}_{T})-f^{\star}\}$, where $D$ measures the heterogeneity between the local functions, $\{\bar{x}_{T}\}$ is a sequence of ergodic server iterates, $f^{\star}$ is the optimal value of Problem (1), and $\tau_{\max}$ is the maximum client delay. Unlike [3, 17] (i) we provide a stronger convergence result for the actual, non-ergodic sequence of server iterates, and (ii) completely eliminate the dependency on function heterogeneity and the maximum delay. Although we do not obtain a scaling of $\sigma^{2}$ with $1/n$ like in [3, 17], this is because we consider at most a single local stochastic gradient step at each iteration, unlike [3, 17] that consider only aggregated gradient steps at the server as iterations. Finally, we note that when the true gradients are available ($\sigma=0$ in Assumption 1), AREA converges to the optimal solution of Problem (1) exponentially fast with the number of iterations $k$ (for empirical results on a toy quadratic problem, see Appendix A.2.1).
@@
In the setting of Example 4, if the client update frequencies $\lambda_{i}$ for $i\in[n]$ are known, one can numerically solve for the aggregation frequency $\lambda_{s}$ which maximizes the convergence rate (i.e., minimizes $\rho(A)$). Moreover, the average time between two consecutive iterations of AREA becomes $1/\bar{\lambda}$, which can be significantly shorter than the time $1/\min_{i}\lambda_{i}$ elapsed between two iterations of synchronous SGD (i.e., waiting for the straggler at each iteration).
We conclude this section with a result on the convergence of AREA when the functions $f_{i}(\cdot)$ of Problem (1) are convex and nonsmooth. In this version of AREA, clients have access only to stochastic approximations of the subgradients of their local functions. Hence, even in the absence of stochastic noise ($\sigma=0$ in Assumption 1), a diminishing stepsize sequence is required to achieve convergence to the set of optimal solutions of Problem (1).
@@
In addition, the sequence of server iterates $\{x_{k}\}$ of AREA converges to a solution of Problem (1) with probability $1$.
Theorem 10 establishes the almost sure convergence of the iterates of AREA to an optimal solution of Problem (1), provided the stepsize sequence $\{\alpha_{k}\}$ is square-summable but not summable. Moreover, for general positive sequence $\{\alpha_{k}\}$ we recover the convergence rate of the centralized stochastic projected subgradient method [7]. For a constant stepsize sequence $\alpha_{k}=\alpha>0$ for $k\in\mathbb{N}_{0}$, taking the total expectation on both sides of Eq. (6) and applying recursively yields $\mathbb{E}\left[f\left(\bar{x}_{s,T}\right)-f^{\star}\right]\leq\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[f(x_{s,k})-f^{\star}]\leq\frac{n}{2}\left(\frac{V_{0}}{\alpha T}+\alpha\hat{\sigma}^{2}\right),$ where $\bar{x}_{s,T}\triangleq\frac{1}{T}\sum_{t=0}^{T-1}x_{s,k}$ and the first inequality holds by Jensen’s inequality. Setting $\alpha=\sqrt{\frac{V_{0}}{\hat{\sigma}^{2}T}}$ and $D_{0}\triangleq\|x_{s,0}-x^{\star}\|$ yields
$\mathbb{E}\left[f\left(\bar{x}_{s,T}\right)-f^{\star}\right]\leq n\hat{\sigma}\sqrt{\frac{V_{0}}{T}}=n\hat{\sigma}D_{0}\sqrt{\frac{p_{s}^{-1}+\frac{2}{n}\sum_{i=1}^{n}p_{i}^{-1}}{T}}.$ (7)
The preceding rate depends only on the average of the inverse probabilities $p_{i}^{-1}$ for $i\in[n]$, rendering AREA less sensitive to excessively fast or slow clients.
To the best of our knowledge, the only known convergence rate for asynchronous SGD under non-iid data and convex, non-smooth functions is $\mathcal{O}\left(\hat{\sigma}D_{0}\sqrt{np_{\max}}/\sqrt{T\sqrt{\bar{p}}}\right)$, where $p_{\max}=\max_{i}p_{i}$ and $\bar{p}=\frac{1}{n}\sum_{i=1}^{n}p_{i}$, as obtained in [14]. The authors of [14] achieve a better dependency on $n$ due to considering each aggregation as an iteration count (we consider both local subgradient steps and aggregations as iterations). However, their rate grows with $p_{\max}$, which may be significantly larger than $\bar{p}$. Moreover, to handle non-iid data, [14] requires a biased sampling scheme where clients are selected according to their importance in the global loss function. In contrast, we allow clients to update the server at their own pace. Finally, in the setting of Example 4, it is possible to obtain a closed form solution for the aggregation frequency $\lambda_{s}$ minimizing $V_{0}$.